OpenStudy (anonymous):

Write the Riemann sum to find the area under the graph of the function f(x) = x4 from x = 5 to x = 7.

jimthompson5910 (jim_thompson5910):

 Hint: \[\Large \text{area under f(x) from a to b} = \lim_{n \to \infty} \sum_{i=1}^{n}f(x_i)*\Delta x\] where \[\Large \Delta x =\frac{b-a}{n}\] \[\Large x_i = a+i*\Delta x\]